OpenStudy (anonymous):

Help i need this answed as soon as possible Complete the square and solve. Show all work. x2 + 4x = 5

jimthompson5910 (jim_thompson5910):

take half of 4, then square it to get (4/2)^2 = 2^2 = 4 so you add 4 to both sides to get this x^2 + 4x + 4 = 5+4 x^2 + 4x + 4 = 9

jimthompson5910 (jim_thompson5910):

the left side is now a perfect square trinomial, which factors to (x+2)^2 so, x^2 + 4x + 4 = 9 (x+2)^2 = 9 I'll let you finish

OpenStudy (anonymous):

thank you :)

jimthompson5910 (jim_thompson5910):

you're welcome